# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 7.1.1 Asker, Fershtman, Jeon, and Pakes (2020): Q-Learning in Dynamic Procurement Auctions
Firms update Q-values using sample averaging:
@@ 7.1.2 Hollenbeck (2019): TD Learning for Merger Analysis with Innovation
The value function update for firm $i$ is100100100This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),101101101I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.102102102The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (62), re-derives best-response strategies from updated values, and repeats. The RL component replaces the exact Bellman iteration step that would be infeasible over the full state space, focusing computational effort on recurrently visited states.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Their contribution is conceptual, bridging the gap between RL-based behavioral models (where agents learn) and the econometric problem (where the researcher estimates structural parameters from data generated by learning agents). Their estimation method itself is standard econometrics rather than RL. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.103103103The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).
@@ 7.2 Auction Equilibria and Mechanism Design
Computing equilibrium bidding strategies in auctions is a long-standing challenge. Closed-form solutions exist only for narrow families of valuation distributions and auction formats, while numerical methods scale poorly with the number of bidders and items. Recent work has applied deep RL to learn equilibrium strategies and optimal mechanisms.
@@ 7.2.1 Brero, Eden, Gerstgrasser, Parkes, and Rheingans-Yoo (2021): RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).104104104A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).105105105In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),106106106Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.
@@
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,107107107The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.108108108DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 7.3 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)109109109DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.
@@ 8 Reinforcement Learning in Games
Computing Nash equilibria is computationally hard. Daskalakis et al. (2009) showed that finding a Nash equilibrium in two-player normal-form games is PPAD-complete, meaning no polynomial-time algorithm is known and none is likely to exist. Extensive-form games with imperfect information inherit this hardness and introduce additional challenges from sequential decision-making and hidden information.110110110PPAD (Polynomial Parity Arguments on Directed graphs) is a complexity class for problems where existence is guaranteed by a parity argument. PPAD-completeness means the problem is as hard as any problem in this class. For comparison, checking whether a strategy profile is a Nash equilibrium is easy (polynomial time); finding one is PPAD-complete. Counterfactual regret minimization (CFR) achieves $O(1/\sqrt{T})$ convergence to $\varepsilon$-Nash equilibrium in two-player zero-sum games (Zinkevich et al., 2008). After $T$ iterations, exploitability decreases as $1/\sqrt{T}$, so reaching exploitability 0.01 requires approximately 10,000 iterations.111111111An $\varepsilon$-Nash equilibrium is a strategy profile where no player can improve her expected payoff by more than $\varepsilon$ through unilateral deviation. As $\varepsilon\to 0$, this converges to an exact Nash equilibrium. This section applies CFR to the durable goods monopoly problem, benchmarking against the theoretical results of Gul et al. (1986) and Ausubel and Deneckere (1989).

### 15 Mar 2026 v2

@@ 9.1.1 Asker, Fershtman, Jeon, and Pakes (2020): Q-Learning in Dynamic Procurement Auctions
Firms update Q-values using sample averaging.
@@ 9.1.2 Hollenbeck (2019): TD Learning for Merger Analysis with Innovation
The value function update for firm $i$ is142142142This stochastic approximation update, introduced by Pakes and McGuire (1994) for dynamic oligopoly computation, is closely related to temporal-difference learning in the RL literature. The original algorithm uses visit-count averaging ($\alpha_{k}=1/k$ where $k$ counts visits to each state) rather than a fixed learning rate.
@@
where $\Pi_{i}(\Omega,\mathbf{a})$ denotes firm $i$’s per-period profit given industry state $\Omega$ and the joint action profile $\mathbf{a}$ (investment, entry/exit, merger decisions),143143143I use $\Pi_{i}$ for firm $i$’s per-period profit to avoid confusion with policy $\pi$. $\gamma$ is the discount factor, and $\Omega^{\prime}$ is the realized next-period state. The algorithm uses $\varepsilon$-decreasing exploration to prevent convergence to locally suboptimal equilibria.
The equilibrium computation follows the Pakes-McGuire iterative scheme.144144144The Pakes-McGuire algorithm (Pakes and McGuire, 1994) computes Markov perfect equilibria by iterating: (1) given current value functions, compute best-response strategies; (2) given strategies, update value functions via simulation. Convergence is not guaranteed but works well in practice. The algorithm simulates long industry histories, updates each firm’s value function via (80), re-derives best-response strategies from updated values, and repeats.
@@
Two related papers merit brief mention. Lomys and Magnolfi (2024) develop structural estimation methods for strategic settings where agents use learning algorithms (specifically regret-minimizing rules) rather than playing a fixed equilibrium. They impose an “asymptotic no-regret” condition as a minimal rationality requirement and derive identification results for payoff parameters. Covarrubias (2022) uses deep RL to study oligopolistic pricing in a New Keynesian framework, representing firms’ pricing policies as neural networks $\pi_{\phi}(a|s)$. The method uncovers multiple equilibria ranging from competitive to collusive pricing.145145145The most prominent example of algorithmic collusion is Calvano et al. (2020), who showed that independent Q-learning agents in a repeated Bertrand pricing game learn to sustain supra-competitive prices and punish deviators without explicit communication. This result and its implications for competition policy are treated in the companion thesis chapter (Rawat, 2026).
@@ 9.2 Auction Equilibria and Mechanism Design
Closed-form equilibrium bidding strategies exist only for narrow families of valuation distributions and auction formats, and numerical methods scale poorly with the number of bidders and items.
@@ 9.2.1 Brero, Eden, Gerstgrasser, Parkes, and Rheingans-Yoo (2021): RL for Sequential Price Mechanisms
Brero et al. (2021) use RL to design optimal sequential price mechanisms (SPMs), a class of indirect auction mechanisms where agents are approached in sequence and offered menus of items at posted prices. SPMs generalize both serial dictatorship and posted-price mechanisms and essentially characterize all strongly obviously strategyproof (SOSP) mechanisms (Pycia and Troyan, 2023).146146146A mechanism is strategyproof if truthful reporting is a dominant strategy. It is obviously strategyproof if this dominance is apparent even to boundedly rational agents; strongly obviously strategyproof (SOSP) adds that dominance holds at every information set (Pycia and Troyan, 2023).
The mechanism design problem is formulated as a partially observable Markov decision process (POMDP).147147147In a POMDP, the agent cannot directly observe the full state. It maintains a belief distribution over possible states, updated via Bayes’ rule as new observations arrive. This captures the mechanism designer’s uncertainty about bidder valuations. The state at round $t$ includes the set of remaining items $\rho_{t}^{\text{items}}\subseteq[m]$, remaining agents $\rho_{t}^{\text{agents}}\subseteq[n]$, the partial allocation $\mathbf{x}_{t}$, and the agents’ (unobserved) valuation functions. The action $a_{t}=(i_{t},\{p_{j}^{t}\}_{j\in\rho_{t-1}^{\text{items}}})$ specifies which agent to visit next and what prices to offer. The observation is the agent’s purchase decision, from which the mechanism can update beliefs about valuations. The reward is the objective function evaluated at the final allocation:
@@
The policy maps from a sufficient statistic of the observation history to actions. Brero et al. (2021) show that this statistic can be represented compactly. The set of remaining items and agents suffices for independent valuations, while the full allocation matrix is needed for correlated valuations. They train the mechanism policy using Proximal Policy Optimization (PPO),148148148Proximal Policy Optimization (Schulman et al., 2017) is a policy gradient algorithm that constrains each update to a trust region, preventing large destabilizing policy changes. It is described in Chapter 2. which handles the discrete action space (agent selection) and continuous action space (price setting) of the POMDP.
@@
The mechanism’s policy maps from state to a price vector over available items. The key technical contribution is making the auction clearing differentiable. In a standard auction, the allocation is an argmax over bids (non-differentiable), and payments depend discontinuously on the allocation. Ravindranath et al. (2024) replace the hard bundle selection with a softmax relaxation,149149149The softmax relaxation replaces the hard allocation $\mathop{\it argmax}_{i}b_{i}$ with a soft allocation $\exp(b_{i}/\tau)/\sum_{j}\exp(b_{j}/\tau)$, which is differentiable and approaches the hard allocation as $\tau\to 0$. enabling analytical gradient computation through the mechanism. The actor loss is the negative expected revenue, and gradients flow directly through the softmax-relaxed allocation and payment rules. The paper explicitly avoids REINFORCE-style estimators, noting that analytical gradients overcome the sample inefficiency and high variance of score-function methods.
@@
Experiments on settings with additive and combinatorial valuations show that the learned mechanisms achieve up to 13% higher revenue than item-wise Myerson optimal auctions, with the largest gains in combinatorial settings where bundle complementarities make item-wise pricing suboptimal. Standard RL baselines (PPO, SAC) are also outperformed, confirming the advantage of exploiting differentiable structure.150150150DDPG was also tested but found to be unstable. DQN is not applicable to continuous price-setting.
@@ 9.3 Macroeconomic Models
Reinforcement learning and deep learning are increasingly used to solve high-dimensional macroeconomic models where grid-based dynamic programming is infeasible. Heterogeneous agent models, in which a distribution of agents with different wealth levels, productivities, or beliefs interact through markets, generate state spaces that grow with the number of agent types and asset positions. Maliar et al. (2021) demonstrate that deep neural networks can approximate policy and value functions in dynamic economic models, achieving accuracy comparable to established projection methods while scaling to problems with dozens of state variables. Fernández-Villaverde et al. (2023) solve a model with financial frictions and an endogenous wealth distribution using deep learning, obtaining global solutions to a problem where perturbation methods fail due to strong nonlinearities. Fernández-Villaverde et al. (2024) provide a systematic treatment of deep learning methods for high-dimensional dynamic programming problems in economics, covering both single-agent and equilibrium settings. Atashbar and Shi (2023) apply deep deterministic policy gradient (DDPG)151151151DDPG extends Q-learning to continuous action spaces by learning a deterministic policy network alongside a Q-function critic. to a real business cycle model, demonstrating that model-free RL can recover near-optimal consumption and investment policies without deriving optimality conditions such as the Euler equation. This is a rapidly growing area; readers seeking a comprehensive methodological treatment are directed to the cited papers and the references therein.
@@ 10 Reinforcement Learning in Games
With multiple agents adapting simultaneously, each agent’s environment includes the others’ changing policies, so the stationary-transition assumption behind single-agent convergence results fails. Shoham et al. (2007) enumerate five desiderata for multi-agent learning: (1) convergence to a stationary strategy in self-play; (2) rationality (best-responding against stationary opponents); (3) equilibrium attainment; (4) safety (guaranteeing at least the Nash-value payoff); (5) social welfare. No existing algorithm satisfies all five in general games.
Two paradigms emerged. Value-based methods generalize the Bellman operator to games, replacing the $\max$ with game-theoretic solution concepts (minimax, Nash), targeting stochastic games with simultaneous moves and observable payoffs. Regret-based methods (CFR) accumulate counterfactual regrets and let the time-averaged strategy converge, targeting extensive-form games with sequential moves and private information. Computing Nash equilibria is PPAD-complete (Daskalakis et al., 2009), so neither approach escapes the fundamental hardness, but both achieve convergence in the game classes they target.
10.1 Stochastic Games and Equilibrium Learning
10.1.1 The Stochastic Game Framework
An $n$-player stochastic game $\Gamma=(n,\mathcal{S},\mathcal{A}_{1},\ldots,\mathcal{A}_{n},P,r_{1},\ldots,r_{n},\gamma)$ consists of a finite state space $\mathcal{S}$; finite action sets $\mathcal{A}_{i}$ for each player $i$; a transition function $P:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\Delta(\mathcal{S})$; reward functions $r_{i}:\mathcal{S}\times\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}\to\mathbb{R}$; and a common discount factor $\gamma\in[0,1)$. At each stage, all players simultaneously choose actions, receive individual rewards, and the game transitions to a new state.152152152Shapley (1953) introduced stochastic games in 1953 for the two-player zero-sum case, proving existence of the value via a contraction argument on the Bellman operator. The general-sum extension to $n$ players is due to Fink (1964) and Takahashi (1964).
A Markov decision process is a stochastic game with $n=1$; a matrix game is a stochastic game with $|\mathcal{S}|=1$. Each player $i$ seeks a policy $\pi_{i}:\mathcal{S}\to\Delta(\mathcal{A}_{i})$ maximizing discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{i}(s_{t},a_{1,t},\ldots,a_{n,t})]$.
Standard Q-learning convergence (Watkins and Dayan, 1992) requires stationary transition and reward dynamics; with multiple learners, this assumption fails. Bowling and Veloso (2002) formalized two properties a learning algorithm should satisfy. It is rational if, when all other players converge to stationary policies, it converges to a best response; it is convergent if, in self-play, it converges to a stationary policy.
If all players use rational, convergent algorithms, the resulting profile is a Nash equilibrium by construction. The challenge is achieving both properties simultaneously.
10.1.2 Minimax-Q Learning
Littman (1994) proposed the first Q-learning algorithm for stochastic games, targeting two-player zero-sum games. The key modification replaces the $\max$ operator in the standard Q-learning backup with a minimax operator. Each agent maintains $Q_{i}(s,a_{i},a_{-i})$ over the joint action space. The update rule is
$Q_{i}(s,a_{i},a_{-i})\leftarrow(1-\alpha)\,Q_{i}(s,a_{i},a_{-i})+\alpha\left[r_{i}+\gamma\,V_{i}(s^{\prime})\right],$ (82)
where the value backup solves a linear program:
$V_{i}(s)=\max_{\pi_{i}\in\Delta(\mathcal{A}_{i})}\min_{a_{-i}\in\mathcal{A}_{-i}}\sum_{a_{i}\in\mathcal{A}_{i}}\pi_{i}(a_{i})\,Q_{i}(s,a_{i},a_{-i}).$ (83)
This is the RL analogue of Shapley’s value iteration for zero-sum stochastic games (Shapley, 1964). The resulting policy $\pi_{i}(s)$ is generally a mixed strategy, since deterministic policies are exploitable in adversarial settings.
Minimax-Q converges to the minimax Q-values under Robbins-Monro learning rates ($\sum_{t}\alpha_{t}=\infty$, $\sum_{t}\alpha_{t}^{2}<\infty$) and infinite exploration of all state-action tuples.153153153The convergence proof extends the contraction argument for standard Q-learning. Because the zero-sum minimax operator is a contraction with modulus $\gamma$ under the $\ell^{\infty}$ norm, the stochastic approximation converges to the fixed point. See Littman (1994) and the general treatment in Szepesvári and Littman (1999). However, the algorithm sacrifices rationality: it plays the equilibrium strategy even against exploitable opponents.154154154In the soccer game of Littman (1994), minimax-Q won 53.7% against a hand-built opponent versus 26.1% for Q-learning. Against an adversarial challenger, Q-learning won 0% (its deterministic policy was fully predictable); minimax-Q won 37.5% through mixed strategies.
10.1.3 Nash-Q Learning
Hu and Wellman (2003) extended the framework to general-sum stochastic games, where players may have aligned, opposed, or mixed incentives. Each agent $i$ maintains a Q-function over the joint action space $Q_{i}(s,a_{1},\ldots,a_{n})$ and updates via
$Q_{i}(s,\mathbf{a})\leftarrow(1-\alpha)\,Q_{i}(s,\mathbf{a})+\alpha\left[r_{i}+\gamma\,\text{Nash}_{i}\bigl(Q_{1}(s^{\prime}),\ldots,Q_{n}(s^{\prime})\bigr)\right],$ (84)
where $\text{Nash}_{i}(\cdot)$ denotes player $i$’s payoff under a Nash equilibrium of the stage game defined by the current Q-values $(Q_{1}(s^{\prime}),\ldots,Q_{n}(s^{\prime}))$. At each backup, the algorithm treats the Q-values as payoff matrices, computes a Nash equilibrium of this matrix game, and uses the equilibrium payoffs for the value estimate.
Theorem 5 (Hu and Wellman (2003)).
Nash-Q converges to Nash Q-values under Robbins-Monro learning rates and infinite exploration, provided every stage game encountered during learning has either (a) a global optimal point (all agents receive their highest payoff at the same joint action) or (b) a unique saddle-point Nash equilibrium.
These conditions are restrictive. When stage games have multiple Nash equilibria, agents may select different equilibria for their backups, causing divergence.155155155In the experiments of Hu and Wellman (2003), a grid game with a unique equilibrium Q-function converged in 100% of trials under Nash-Q versus 20% under independent Q-learning; a game with three equilibrium Q-functions converged in only 68–90% of trials. Nash-Q requires each agent to observe all other agents’ rewards and to maintain Q-values over the joint action space $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$, with storage $O(n|\mathcal{S}|\prod_{i}|\mathcal{A}_{i}|)$, exponential in the number of players. Nash-Q reduces to standard Q-learning in the single-agent case.
10.1.4 The Convergence Problem
Table 18 summarizes the trade-offs. No single algorithm achieves both rationality and convergence in general games.
Table 18: Multi-agent Q-learning algorithms: convergence and information requirements
Algorithm Rational Convergent Game class Information
Indep. Q-learning Yes No Any Own reward
Minimax-Q No Yes Zero-sum Joint actions
Nash-Q Cond. Cond. General-sum All rewards
WoLF-PHC Yes Yes ($2{\times}2$) General-sum Own reward
WoLF-PHC (Win or Learn Fast, Policy Hill-Climbing) of Bowling and Veloso (2002) maintains a policy $\pi_{i}(a|s)$ and a running average policy $\bar{\pi}_{i}(a|s)$, updating Q-values as in standard Q-learning. The policy moves toward the greedy action at a variable rate:
$\delta=\begin{cases}\delta_{l}&\text{if }\sum_{a}\pi_{i}(a|s)\,Q_{i}(s,a)<\sum_{a}\bar{\pi}_{i}(a|s)\,Q_{i}(s,a)\quad\text{(losing)}\\ \delta_{w}&\text{otherwise}\quad\text{(winning)}\end{cases}$ (85)
with $\delta_{l}>\delta_{w}$. When the current policy underperforms the historical average (losing), the agent adapts quickly; when outperforming (winning), it adapts slowly to avoid destabilizing the opponent. Bowling and Veloso (2002) proved that WoLF-IGA (the infinitesimal gradient ascent variant) converges to Nash equilibrium in all two-player, two-action games. The trajectory traces piecewise ellipses around the equilibrium, shrinking by a factor of $\ell^{4}<1$ per orbit, where $\ell=\sqrt{\delta_{w}/\delta_{l}}$. WoLF-PHC requires only own-reward observations, the same information as independent Q-learning.
Two further algorithms deserve mention. Friend-or-Foe Q-learning (Littman, 2001) decomposes agents as cooperative (friend) or adversarial (foe), using $\max$ for friends and minimax for foes; it always converges but requires knowing the relationship type a priori.156156156Correlated-Q learning (Greenwald and Hall, 2003) generalizes both Nash-Q and Minimax-Q by using correlated equilibrium, a probability distribution over joint actions enforced by a correlating device. The set of correlated equilibria contains all Nash equilibria. Correlated-Q converges under conditions analogous to Nash-Q. The evolutionary perspective of Börgers and Sarin (1997) provides a deeper lens: reinforcement learning dynamics in matrix games converge to the replicator equation from evolutionary game theory. The cycling of Q-learning in games such as matching pennies is structurally identical to the cycling of replicator dynamics in Rock-Paper-Scissors games.
10.1.5 Simulation Study: Cournot and Bertrand Duopoly
Two canonical games from industrial organization serve as benchmarks. In Cournot duopoly, two firms choose quantities $q_{i}\in\{0,1,\ldots,9\}$ with inverse demand P(Q)=10QP(Q)=10-Q and marginal cost $c=1$; the unique Nash equilibrium is $q^{*}=3$ with profit $\pi^{*}=9$. In Bertrand duopoly with differentiated products, two firms choose prices $p_{i}\in\{0,1,\ldots,9\}$ with demand $d_{i}=10-2p_{i}+p_{j}$ and marginal cost $c=1$; the continuous Nash equilibrium is $p^{*}\approx 4.33$, which discretizes to $p^{*}=4$ with profit $\pi^{*}=18$. Both games have unique Nash equilibria in pure strategies.157157157Each configuration runs for 50,000 iterations across 20 seeds. Three algorithms compete: independent Q-learning (IQL), Nash-Q, and WoLF-PHC.
Table 19: Convergence to Nash equilibrium in Cournot and Bertrand duopoly
Game Algorithm Action Profit $|a-a^{*}|$ Conv. iter
Cournot IQL $2.95\pm 0.05$ $9.1\pm 0.0$ $0.05$ 1,000
Nash-Q $2.89\pm 0.33$ $8.8\pm 1.3$ $0.17$ 1,000
WoLF-PHC $3.00\pm 0.00$ $9.0\pm 0.0$ $0.00$ 1,000
Bertrand IQL $4.00\pm 0.00$ $18.0\pm 0.0$ $0.33$ 1,000
Nash-Q $4.00\pm 0.00$ $18.0\pm 0.0$ $0.33$ 1,000
WoLF-PHC $3.95\pm 0.05$ $17.8\pm 0.2$ $0.38$ 1,000
Notes: Action and Profit report mean $\pm$ standard error across 20 seeds over the final 5,000 iterations. $|a-a^{*}|$ is the mean distance from Nash. Conv. iter is the first iteration where the smoothed average action enters a $0.5$-neighborhood of Nash.
Table 19 and Figure 17 report the results. All three algorithms converge to Nash in both games within the first 5,000 iterations. IQL, which lacks any game-theoretic computation, matches the game-aware methods in these well-structured games with unique pure-strategy equilibria. The advantage of Nash-Q and WoLF-PHC emerges in games requiring mixed strategies, where IQL’s deterministic policy limit prevents convergence.
Figure 17: Convergence of expected actions to Nash equilibrium. Left: Cournot duopoly. Right: Bertrand duopoly. Smoothed over 1,000-iteration windows, averaged across 20 seeds.